# Black holes surrounded by generic matter distributions: polar perturbations and energy flux

### v1

@@ II Background and Density Profiles
Following the Einstein cluster prescription Einstein (1939); Geralico et al. (2012), the properties of the environment are described by an anisotropic stress-energy tensor $(T^{\text{env}})^{\mu}{{}_{\nu}}=\text{diag}(-\rho,0,P_{t},P_{t})$, where $\rho(r)$ and $P_{t}(r)$ are the density and the tangential pressure of the matter distribution. The continuity equation determines the mass profile through the relation $m^{\prime}(r)=4\pi\rho(r)$. The metric functions $a(r)$ and $P_{t}(r)$ are found from the $rr$ component of Eqs. (2) and from the Bianchi identities, respectively:
@@ II.1 Dark matter density profiles
Figure 1: Mass profiles for some representative DM configurations considered in this work. The NFW model is truncated at the cutoff radius $r_{c}=5a_{0}$. The Hernquist, NFW and Einasto profiles are obtained by the “cutoff factor” recipe $\rho(r)\rightarrow(1-2M_{\text{BH}}/r)\rho(r)$. The “DM spike” profile is obtained with the fully relativistic calculation of the adiabatic growth of a BH within a Hernquist profile from Refs. Sadeghian et al. (2013); Speeney et al. (2022), as described in the main text. Going from the top to the bottom panels, the DM halo mass grows relative to the central BH mass ($M_{\text{Halo}}/M_{\text{BH}}=1,\,10,\,100$), while the compactness of the halo decreases going from the left to the right panels ($M_{\text{Halo}}/a_{0}=0.1,\,0.01,\,0.001$). The vertical lines correspond to a representative orbital radius $r_{p}=7.9456M_{\text{BH}}$ that will be used in Table 1 below to compare GW energy fluxes.
@@
The DM density profile arising from the (adiabatic) accretion growth of a nonrotating BH sitting at the core is expected to develop an overdense cusp followed by a sharp cutoff close to the horizon Gondolo and Silk (1999). A general relativistic analysis predicts that the density falls off at $r=4M_{\text{BH}}$ Sadeghian et al. (2013). In our numerical profiles, we mimic this behavior by considering two different approaches. The first approach consists of following the strategy adopted in Ref. Figueiredo et al. (2023), i.e., we multiply the density of the DM models discussed above by a “cutoff factor” such that $\rho(r)\rightarrow(1-2M_{\text{BH}}/r)\rho(r)$. In the second approach we consider the fully relativistic “DM spike” models investigated by some of us in previous work Speeney et al. (2022). In that work, following Ref. Sadeghian et al. (2013), we derived semianalytic fits valid for $r\lesssim a_{0}$ for spiked profiles based on the Hernquist and NFW distributions, and used these fits to assess the impact of the overdensities on the GW emission of binary sources.
From now on we will focus on the Hernquist subclass of the DM spikes, i.e., we consider the adiabatic growth of a spike starting from Eq. (7) with $(\alpha,\beta,\gamma)=(1,4,1)$. This is because the Hernquist and NFW DM spikes are very similar at small values of $r$ (see Fig. 1) and they mostly differ at large values of $r$, which are not of interest for our present goals.
@@ III Numerical Procedure
We first make a specific choice of $\rho(r)$, and then we solve for the background metric component $a(r)$ and for the mass function $m(r)$. This is accomplished by numerically integrating $m^{\prime}(r)=4\pi r^{2}\rho(r)$ from $r=2M_{\text{BH}}$ to the cutoff radius $r_{\infty}=10^{11}M_{\text{BH}}$, which corresponds to our numerical approximation of spatial infinity. We then plug the numerical solution for $m(r)$ into Eq. (3) to solve for $a(r)$ by integrating “backwards” from $r_{\infty}$, with the following boundary condition:
@@
where $\{K,\,H_{0},\,H_{1}\}$ and $\{W,\,\delta\rho\}$ are the perturbation variables related to the metric and to the fluid, respectively. The components of the matrix $\boldsymbol{\alpha}$ and of the source vector $\vec{S}$ are given explicitly in Ref. Cardoso et al. (2022b). We require the gravitational perturbation variables to satisfy ingoing- and outgoing-wave boundary conditions at the horizon and at spatial infinity, respectively. Following Ref. Cardoso et al. (2022b) we assume, for simplicity, that the fluid variables vanish at the two boundaries, i.e. $W(2M_{\text{BH}})=\delta\rho(2M_{\text{BH}})=W(r_{\infty})=\delta\rho(r_{\infty})=0$. The gravitational perturbations at the horizon can be expanded as follows:
@@
$H_{0}^{(\text{in})}(r)$ $=e^{-i\omega r*}\sum_{i=0}^{N_{\text{in}}}H_{0,H}^{(i)}(r-2M_{\text{BH}})^{i}\ ,$
$H_{1}^{(\text{in})}(r)$ $=e^{-i\omega r*}\sum_{i=0}^{N_{\text{in}}}H_{1,H}^{(i)}(r-2M_{\text{BH}})^{i}\ ,$
@@
$2~{}2$ 1.7068e-4 1 100 2.4341e-4   1.7038e-4 1.4133e-4   1.7037e-4 2.1997e-4   1.7036e-4 1.4396e-4   1.7036e-4 8.0265e-4   1.6972e-4 5.8136e-4   1.6833e-4 1.8033e-4   1.6997e-4 1.6533e-4   1.7037e-4
$3~{}1$ 2.2443e-9 1 100 5.8115e-9   2.1823e-9 2.3212e-9   2.2010e-9 4.2975e-10   2.1788e-9 2.1489e-9   2.2013e-9 2.3454e-6   2.4672e-9 7.1498e-8   2.1772e-9 5.4913e-9   2.4933e-9 3.1469e-9   2.2568e-9
$3~{}3$ 2.5488e-5 1 100 3.5564e-5   2.5429e-5 2.1075e-5   2.5448e-5 3.1495e-5   2.5429e-5 2.1446e-5   2.5434e-5 1.4468e-4   2.5301e-5 7.6403e-6   2.5131e-5 2.5779e-5   2.5323e-5 2.4470e-5   2.5415e-5
$4~{}2$ 2.5855e-9 1 100 3.1319e-8   2.4966e-9 1.1507e-9   2.5389e-9 2.5378e-8   2.5028e-9 1.5796e-9   2.5384e-9 1.9289e-7   2.0099e-9 3.3474e-9   2.5045e-9 4.6899e-9   2.0384e-9 8.3514e-10   2.4449e-9
$4~{}4$ 4.7327e-6 1 100 6.0451e-6   4.7169e-6 3.8716e-6   4.7200e-6 5.2628e-6   4.7177e-6 3.9592e-6   4.7204e-6 3.6665e-5   4.6776e-6 1.2052e-6   4.6639e-6 4.3577e-6   4.6794e-6 4.4617e-6   4.7165e-6
Table 1: Selected polar multipolar components $\dot{E}^{\infty}_{(\ell,m)}$ of the energy flux at spatial infinity in units of $q^{2}=m_{p}^{2}/M_{\text{BH}}^{2}$ for a particle orbiting at $r_{p}=7.9456M_{\text{BH}}$. In each case, we set the width of the Gaussian to $\sigma=1/50$. For nonvacuum fluxes, the four entries in each cell span two values of the DM halo mass and compactness: $M_{\text{Halo}}/M_{\text{BH}}=1,\,100$, $M_{\text{Halo}}/a_{0}=0.1,\,0.001$.
@@
$H_{0}^{(\text{out})}(r)$ $=e^{i\omega r*}\sum_{i=0}^{N_{\text{out}}}\frac{H_{0,\infty}^{(i)}}{r^{i}}\ ,$
$H_{1}^{(\text{out})}(r)$ $=e^{i\omega r*}\sum_{i=0}^{N_{\text{out}}}\frac{H_{1,\infty}^{(i)}}{r^{i}}\ ,$
@@ IV Results
The solid lines in the top-left panel of Fig. 3 show the relative difference between the $\ell=m=2$ multipolar contribution to the energy flux in the frequency domain for the Hernquist distribution and the corresponding vacuum result for different DM halos with fixed compactness $M_{\textnormal{Halo}}/a_{0}=0.01$. The plot spans values of the orbital frequencies such that the circular orbit radius of the particle varies between $r_{p}=50M_{\text{BH}}$ and $r_{p}=6M_{\text{BH}}$. Results are in excellent agreement with those in Fig. 3 of Ref. Cardoso et al. (2022b), which made use of closed-form analytic expressions for the metric functions $m(r)$ and $a(r)$. In fact, the energy fluxes shown here are smoother compared to those of Ref. Cardoso et al. (2022b), mainly because of the larger value of $r_{\text{out}}$ considered in this paper to increase numerical accuracy.
@@ V Conclusions
In Fig. 2 we focus on the fiducial “DM spike” model, computed from a fully relativistic calculation of the adiabatic growth of a BH within the Hernquist profile Sadeghian et al. (2013); Speeney et al. (2022). The quadrupolar energy flux decreases as the DM halo become less compact at fixed values of $M_{\text{Halo}}/M_{\text{BH}}$, because the strong-field dynamics of the particle is dominated by the BH when the DM profile is more dilute (i.e., when the DM density close to the BH decreases). The energy flux is smaller for large values of $M_{\text{Halo}}/M_{\text{BH}}$ at fixed compactness for a similar reason: more dilute DM profiles have less DM within the particle orbit. Realistic DM profiles are probably in the regime $M_{\text{Halo}}/a_{0}\ll 1$, and we plan to explore more carefully this “dilute DM” limit in future work.
@@
First of all, it is important to accurately model dynamical friction, which may accumulate large dephasings over the $\mathcal{O}(10^{5}-10^{6})$ orbital cycles typical of LISA EMRIs Eda et al. (2015); Cardoso and Maselli (2020); Speeney et al. (2022); Coogan et al. (2022); Cardoso et al. (2019); Traykova et al. (2021, 2023). Moreover, our model does not properly take into account matter fluxes. While EMRIs can be correctly described by stationary configurations with matter profiles that vanish at the horizon and at spatial infinity, intermediate-mass binaries will be sensitive to feedback effects. This backreaction will affect the DM profile during the evolution of the orbit Kavanagh et al. (2020); Coogan et al. (2022); Nichols et al. (2023). A proper treatment of backreaction requires the inclusion of higher-order terms in the perturbative expansion. If simple estimates similar to Eq. (20) hold at very small orbital frequencies, DM halos could produce effects of phenomenological relevance for pulsar timing arrays Ghoshal and Strumia (2023).
On the data analysis side, prior works have proposed to use EMRI observations to distinguish various environmental signatures Cole et al. (2022, 2023); Barausse et al. (2014, 2015); Speeney et al. (2022); Vicente and Cardoso (2022). The majority of these works uses a combination of fully relativistic and PN estimates of environmental effects, but so far there is no consistent description of the evolution of the BH spacetime embedded in DM during the inspiral. Our framework can accommodate relativistic effects in the generation and propagation of GWs. In the future, we plan to develop waveform models based on these flux calculations and to use them in a Bayesian framework to estimate the detectability of DM overdensities.

### v2

@@ II Background and Density Profiles
Following the Einstein cluster prescription Einstein (1939); Geralico et al. (2012), the properties of the environment are described by an anisotropic stress-energy tensor $(T^{\text{env}})^{\mu}{{}_{\nu}}=\text{diag}(-\rho,0,P_{t},P_{t})$, where $\rho(r)$ and $P_{t}(r)$ are the density and the tangential pressure of the matter distribution. The continuity equation determines the mass profile through the relation $m^{\prime}(r)=4\pi r^{2}\rho(r)$. The metric functions $a(r)$ and $P_{t}(r)$ are found from the $rr$ component of Eqs. (2) and from the Bianchi identities, respectively:
@@ II.1 Dark matter density profiles
Figure 1: Mass profiles for some representative DM configurations considered in this work. The NFW model is truncated at the cutoff radius $r_{c}=5a_{0}$. The Hernquist, NFW and Einasto profiles are obtained by the “cutoff factor” recipe $\rho(r)\rightarrow(1-4M_{\text{BH}}/r)\rho(r)$. The “DM spike” profile is obtained with the fully relativistic calculation of the adiabatic growth of a BH within a Hernquist profile from Refs. Sadeghian et al. (2013); Speeney et al. (2022), as described in the main text. Going from the top to the bottom panels, the DM halo mass grows relative to the central BH mass ($M_{\text{Halo}}/M_{\text{BH}}=1,\,10,\,100$), while the compactness of the halo decreases going from the left to the right panels ($M_{\text{Halo}}/a_{0}=0.1,\,0.01,\,0.001$). The vertical lines correspond to a representative orbital radius $r_{p}=7.9456M_{\text{BH}}$ that will be used in Table 1 below to compare GW energy fluxes.
@@
The DM density profile arising from the (adiabatic) accretion growth of a nonrotating BH sitting at the core is expected to develop an overdense cusp followed by a sharp cutoff close to the horizon Gondolo and Silk (1999). A general relativistic analysis predicts that the density falls off at $r=4M_{\text{BH}}$ Sadeghian et al. (2013). In our numerical profiles, we mimic this behavior by considering two approaches. The first approach consists of following a strategy similar to that adopted in Ref. Figueiredo et al. (2023), i.e., we multiply the density of the DM models discussed above by a “cutoff factor” such that $\rho(r)\rightarrow(1-4M_{\text{BH}}/r)\rho(r)$. The prefactor of $(1-4M_{\text{BH}}/r)$, rather than the prefactor of $(1-2M_{\text{BH}}/r)$ used in in Ref. Figueiredo et al. (2023), is chosen to avoid potential problems with the tangential pressure which arise in the region $2M_{\text{BH}}<r<4M_{\text{BH}}$ Datta (2023): with the present choice, the ratio $P_{t}(r)/\rho(r)$ remains finite and smaller than unity for all $r$. In the second approach we consider the fully relativistic “DM spike” models investigated by some of us in previous work Speeney et al. (2022). In that work, following Ref. Sadeghian et al. (2013), we derived semianalytic fits valid for $r\lesssim a_{0}$ for spiked profiles based on the Hernquist and NFW distributions, and used these fits to assess the impact of the overdensities on the GW emission of binary sources.
From now on we will focus on the Hernquist subclass of the DM spikes, i.e., we consider the adiabatic growth of a spike starting from Eq. (7) with $(\alpha,\beta,\gamma)=(1,4,1)$. This is because the Hernquist and NFW DM spikes are very similar at small values of $r$ (see Fig. 1).
@@
Note that in all the cases we consider, the DM halo density and compactness are quite high compared to astrophysical galactic halos. Observations of SgrA at the Galactic centre from the GRAVITY collaboration show that the extended mass in the region $r\lesssim 10^{6}M_{\text{BH}}$ is very small, with the total mass function dominated by the central $4.3\times 10^{6}M_{\odot}$ BH Abuter et al. (2022). The distributions considered in this work have masses $\sim(1-10)M_{\text{BH}}$ within radii $r<10^{4}M_{\text{BH}}$, and they were mainly chosen to provide a proof of concept that our computational pipeline can be used to study relativistic environmental effects. These distribution may still be representative of DM halos in galaxies smaller than the Milky Way, or of subhalos Springel et al. (2008).
@@ III Numerical Procedure
We first make a specific choice of $\rho(r)$, and then we solve for the background metric component $a(r)$ and for the mass function $m(r)$. In lieu of analytic or approximate solutions as found in Shen et al. (2023); Konoplya and Zhidenko (2022), we choose to numerically solve for the metric components by numerically integrating $m^{\prime}(r)=4\pi r^{2}\rho(r)$ from $r=2M_{\text{BH}}$ to the cutoff radius $r_{\infty}=10^{11}M_{\text{BH}}$, which corresponds to our numerical approximation of spatial infinity. We then plug the numerical solution for $m(r)$ into Eq. (3) to solve for $a(r)$ by integrating “backwards” from $r_{\infty}$, with the following boundary condition:
@@
where $\{K,\,H_{0},\,H_{1}\}$ and $\{W,\,\delta\rho\}$ are the perturbation variables related to the metric and to the fluid, respectively. The components of the matrix $\boldsymbol{\alpha}$ and of the source vector $\vec{S}$ are given explicitly in Ref. Cardoso et al. (2022b). We require the gravitational perturbation variables to satisfy ingoing- and outgoing-wave boundary conditions at the horizon and at our numerical infinity, respectively. Following Ref. Cardoso et al. (2022b) we assume, for simplicity, that the fluid variables vanish at the two boundaries, i.e. $W(2M_{\text{BH}})=\delta\rho(2M_{\text{BH}})=W(r_{\infty})=\delta\rho(r_{\infty})=0$. The gravitational perturbations at the horizon can be expanded as follows:
@@
$H_{0}^{(\text{in})}(r)$ $=e^{-i\omega r*}\sum_{i=0}^{N_{\text{in}}}H_{0,H}^{(i)}(r-2M_{\text{BH}})^{i-1}\ ,$
$H_{1}^{(\text{in})}(r)$ $=e^{-i\omega r*}\sum_{i=0}^{N_{\text{in}}}H_{1,H}^{(i)}(r-2M_{\text{BH}})^{i-1}\ ,$
@@
$2~{}2$ 1.7068e-4 1 100 1.9727e-4   1.7029e-4 1.4045e-4   1.7035e-4 1.8620e-4   1.7035e-4 1.4352e-4   1.7034e-4 2.9443e-4   1.6915e-4 4.7683e-4   1.6831e-4 1.8033e-4   1.6997e-4 1.6533e-4   1.7037e-4
$3~{}1$ 2.2443e-9 1 100 5.9746e-9   2.1793e-9 2.1349e-9   2.1742e-9 9.0170e-10   2.1774e-9 2.0365e-9   2.1746e-9 5.8263e-6   2.3559e-9 3.0898e-8   2.1500e-9 5.4913e-9   2.4933e-9 3.1469e-9   2.2568e-9
$3~{}3$ 2.5488e-5 1 100 2.8080e-5   2.5421e-5 2.0893e-5   2.5441e-5 2.6209e-5   2.5429e-5 2.1384e-5   2.5428e-5 2.8465e-5   2.5217e-5 6.2005e-6   2.5124e-5 2.5779e-5   2.5323e-5 2.4470e-5   2.5415e-5
$4~{}2$ 2.5855e-9 1 100 1.5745e-8   2.5000e-9 1.4164e-9   2.5099e-9 1.3136e-8   2.5050e-9 1.7439e-9   2.5095e-9 8.3112e-10   2.1609e-9 2.1759e-9   2.4770e-9 4.6899e-9   2.0384e-9 8.3514e-10   2.4449e-9
$4~{}4$ 4.7327e-6 1 100 4.6286e-6   4.7162e-6 3.8484e-6   4.7172e-6 4.3036e-6   4.7176e-6 3.9524e-6   4.7177e-6 5.7701e-6   4.6656e-6 9.4731e-7   4.6612e-6 4.3577e-6   4.6794e-6 4.4617e-6   4.7165e-6
Table 1: Selected polar multipolar components $\dot{E}^{\infty}_{(\ell,m)}$ of the energy flux at spatial infinity in units of $q^{2}=m_{p}^{2}/M_{\text{BH}}^{2}$ for a particle orbiting at $r_{p}=7.9456M_{\text{BH}}$. In each case, we set the width of the Gaussian to $\sigma=1/200$. For nonvacuum fluxes, the four entries in each cell span two values of the DM halo mass and compactness: $M_{\text{Halo}}/M_{\text{BH}}=1,\,100$, $M_{\text{Halo}}/a_{0}=0.1,\,0.001$.
@@
$H_{0}^{(\text{out})}(r)$ $=e^{i\omega r*}\sum_{i=0}^{N_{\text{out}}}\frac{H_{0,\infty}^{(i)}}{r^{i-1}}\ ,$
$H_{1}^{(\text{out})}(r)$ $=e^{i\omega r*}\sum_{i=0}^{N_{\text{out}}}\frac{H_{1,\infty}^{(i)}}{r^{i-1}}\ ,$
@@ IV Results
The solid lines in the top-left panel of Fig. 3 show the relative difference between the $\ell=m=2$ multipolar contribution to the energy flux in the frequency domain for the Hernquist distribution and the corresponding vacuum result for different DM halos with fixed compactness $M_{\textnormal{Halo}}/a_{0}=0.01$. The plot spans values of the orbital frequencies such that the circular orbit radius of the particle varies between $r_{p}=50M_{\text{BH}}$ and $r_{p}=6M_{\text{BH}}$. We have checked that when we use a cutoff factor of $(1-2M_{\text{BH}}/r)$ our results are in excellent agreement with those in Fig. 3 of Ref. Cardoso et al. (2022b), which made use of closed-form analytic expressions for the metric functions $m(r)$ and $a(r)$. In fact, the energy fluxes shown here are smoother compared to those of Ref. Cardoso et al. (2022b), mainly because of the larger value of $r_{\text{out}}$ considered in this paper to increase numerical accuracy.
@@ V Conclusions
In Fig. 2 we focus on the fiducial “DM spike” model, computed from a fully relativistic calculation of the adiabatic growth of a BH within the Hernquist profile Sadeghian et al. (2013); Speeney et al. (2022). The quadrupolar energy flux decreases as the DM halo become less compact at fixed values of $M_{\text{Halo}}/M_{\text{BH}}$, because the strong-field dynamics of the particle is dominated by the BH when the DM profile is more dilute (i.e., when the DM density close to the BH decreases). The energy flux is smaller for large values of $M_{\text{Halo}}/M_{\text{BH}}$ at fixed compactness for a similar reason: more dilute DM profiles have less DM within the particle orbit. Realistic DM profiles are probably in the regime $M_{\text{Halo}}/a_{0}\ll 1$, as is the case for a Milky Way-like galaxy. We plan to explore more carefully this “dilute DM” limit in future work.
@@
First of all, it is important to accurately model dynamical friction, which may accumulate large dephasings over the $\mathcal{O}(10^{5}-10^{6})$ orbital cycles typical of LISA EMRIs Eda et al. (2015); Cardoso and Maselli (2020); Speeney et al. (2022); Coogan et al. (2022); Cardoso et al. (2019); Traykova et al. (2021, 2023). Moreover, our model does not properly take into account matter fluxes. While EMRIs can be correctly described by stationary configurations with matter profiles that vanish at the horizon and at spatial infinity, intermediate-mass binaries will be sensitive to feedback effects. This backreaction will affect the DM profile during the evolution of the orbit Kavanagh et al. (2020); Coogan et al. (2022); Nichols et al. (2023). A proper treatment of backreaction requires the inclusion of higher-order terms in the perturbative expansion. If simple estimates similar to Eq. (20) hold at very small orbital frequencies, DM halos could produce effects of phenomenological relevance for pulsar timing arrays Ghoshal and Strumia (2023); Aghaie et al. (2023).
On the data analysis side, prior works have proposed to use EMRI observations to distinguish various environmental signatures Cole et al. (2022, 2023); Barausse et al. (2014, 2015); Speeney et al. (2022); Vicente and Cardoso (2022); Rahman et al. (2023). The majority of these works use a combination of fully relativistic and PN estimates of environmental effects, but so far there is no consistent description of the evolution of the BH spacetime embedded in DM during the inspiral. Our framework can accommodate relativistic effects in the generation and propagation of GWs. In the future, we plan to develop waveform models based on these flux calculations and to use them in a Bayesian framework to estimate the detectability of DM overdensities.